The number of deer in a wildlife preserve is given by the function D(t)=184(1.1)^0.5t, where t is time in years. How long does it take the deer population to triple?
Answer by stanbon(75887)

The number of deer in a wildlife preserve is given by the function D(t)=184(1.1)^0.5t, where t is time in years. How long does it take the deer population to triple?
-------------
3(184)=184(1.1)^0.5t
(1.1)^0.5t=3
Take the log of both sides to ge:
0.05t(log11)=log3
0.05t=0.45815691..
t=9.16313... yrs
Cheers,
Stan H.